OpenStudy (anonymous):

The sum of the four angle measures of any convex quadrilateral is 360 degrees. Suppose that a convex quadrilateral has angle measures of 60, 8y+3, and 6y-7 degrees. Write an expression for the degree measure of the fourth angle. @uri @Becki @ChasingMaggie @kah.x @sangya21

OpenStudy (anonymous):

60+8y+3+6y-7+x=360 where x=the 4th angle of the quadrilateral

OpenStudy (anonymous):

So, the fourth angle would be 360 minus the other three angles, do you agree?

OpenStudy (anonymous):

okay so the degree measure of the 4th angle is just 360 as the answer

OpenStudy (anonymous):

No. 360 is the number of degrees in all four angles total. They gave you 3 of them and told you the total (360).

OpenStudy (anonymous):

okay so how do i find the answrr

OpenStudy (anonymous):

x=360 -(all the other angles)

OpenStudy (anonymous):

See my first response. I did everything but write it in equation form for you.

OpenStudy (anonymous):

all the other angles are 60, 8y+3 and 6y-7?

OpenStudy (anonymous):

im confused what im subtravting?

OpenStudy (anonymous):

@sangya21 what are the other angles?

OpenStudy (anonymous):

x = 360-(60+8y+3+6y-7)

OpenStudy (anonymous):

oh okay thank you.

OpenStudy (anonymous):

can you also help with this one Latoya has a job transporting soft drinks by truck. Her truck is filled with cans that weigh 14 ounces each and bottles that weigh 70 ounces each. There is a combined total of 770 cans and bottles in her truck. Let x be the number of 14 -ounce cans in her truck. Write an expression for the combined total weight (in ounces) of the cans and bottles in her truck.

OpenStudy (anonymous):

@sangya21 @kah.x

OpenStudy (anonymous):

14x +70(770-x)

OpenStudy (anonymous):

Two trains leave stations 588 miles apart at the same time and travel toward each other. One train travels at 115 miles per hour while the other travels at 95 miles per hour. How long will it take for the two trains to meet?

OpenStudy (anonymous):

@kah.x @sangya21

OpenStudy (anonymous):

this question is the one with which I always have trouble.

OpenStudy (anonymous):

yeah me too well all "word problems"

OpenStudy (anonymous):

@Becki do u know?